OpenStudy (anonymous):

draw the graph that matches this equation. -6/7x - 1/2y = 3/7

OpenStudy (anonymous):

rearrange the equation and multiplying both side with factor of 14 -2x - 7y = 6 7y = -2x - 6 y = -2/7 x - 6/7 so the graph will be a straight line with gradient -2/7 and y intercept of -6/7